answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability that an offspring will be male?

For humans, approximately 0.52

For humans, approximately 0.52

For humans, approximately 0.52

For humans, approximately 0.52

What is a queuing system theory?

It is a study of queues. A typical example is one that may be found in banks. Customers arrive at intervals that are determined by some probability distribution function. There are then two possible queueing scenarios: one in which there is a single queue which feeds into several cashiers or a system where there are multiple queues: one for each cashier. When a customer reaches a cashier, it takes the cashier an amount of time to serve him or her. This time also has a probability distribution function - different from the one governing arrival times.

The theory studies the optimum queueing scenario, time spent by customers in queues rather than being served, the optimum number of cashiers. The bank must find the best trade-off between the cost of employing more cashiers and the irritation of their customers.

What is the human probability that a offspring will be a female?

it depends on how much sexual intercourse you have. it is up to you.

What is it called when twin resulting from a fertilized egg splits into two?

Identicle twins are formed from a fertilized embryo that splits into two. Fraternal twins are formed when the woman released to eggs in a cycle and they both get fertilized.

What is the probability that a team will win the a game if the reported odds are a to b?

It depends on whether the odds are in favour or against.

Generally, the odds refer to a wins versus b losses. That is a wins out of a total number of a+b outcomes and so a probability of winning = a/(a+b).

What is the normal distribution of 65 seconds and standard deviation 0.8?

If X is Normally distributed with mean 65 seconds and sd = 0.8 seconds, then Z = (X - 65)/0.8 has a Standard Normal distribution; that is, Z has a N(0, 1) distribution.

The cumulative distribution for Z is easily available - on the net and in any basiic book on statistics. To get to the cumulative dirtribution function of X all you need is to use the transformation X = 0.8*Z + 65.

What is the probability of rolling a seven with two dice given that at least one of them is a four?

If one die is constrained to be a 4, then the probabilities rest with the other die. A 7 in total, then, means that the other die is a 3 or higher. The probability of a three or higher is 4 in 6, or 2 in 3.

What is the probability of flipping four coin's to get a head?

The probability of NOT getting heads is (1/2)4=1/16

Therefore the probability of getting heads is 1-1/16=15/16

What is the probability of a card being a heart?

One in four for any random draw. In mathematical terms 1:4.

When can a binomial situation be approxiamted with a normal distribution?

The binomial distribution can be approximated with a normal distribution when np > 5 and np(1-p) > 5 where p is the proportion (probability) of success of an event and n is the total number of independent trials.

Ten students are playing in a chess tournament If each student plays every other student once how many total games are played?

If you choose one person, then consider your options for a second, you have nine people for them to play against. Then, considering the second person, excluding the match already played with the first person, you have eight possible matches. Following this all the way, the number of unique matches are:

9 + 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 = 45

How many ways can you choose 5 hearts from a 52-card deck?

there are 13 hearts in all so you would use a combination permutation because order does not matter so 13C5 = 1287

What is the Role of Independence in the topic of probabilities and is there an example of it?

There is a wonderful and brief explanation at the link. One thought: Without the concept of independence, the accurate probability that two events will occur together would be a problem. You need to know if the events are dependent on one another in some way. If I roll two fair dice, what is the probability that I will roll two sixes? I know that the events (the results I get from the two dice) are independent of one another. So the probability of their happening together is the product of the probabilities that they will happen independently.

Is the Wonderlic Test used outside of pro sports?

Check out the Wonderlic website - www.wonderlic.com Wonderlic tests are used across the board by companies, large and small, during the pre-employment process. Wonderlic assessments are a proven way for a company to cut down on turnover costs, and hire betting fitting employees.

Answer The "Wonderlic" used to test major league football players is more specifically known as the Wonderlic Personnel Test, or WPT. The WPT is a brief test of general intelligence, the scores of which can easily be converted to IQ scores using a formula.

Years of experience and social sciences research have shown that general intelligence, increasingly referred to as "g" by researchers, is the best single predictor of success in a job or in school (this does not justify using only a single test to make important hiring or acceptance decisions; see below). The predictive usefulness of general intelligence is the fundamental rationale for using the WPT as a pre-employment test in a great variety of occupations, ranging from pro football player to bank teller, police officer, engineering technician, or real estate sales agent.

Although useful, WPT is not the most accurate or comprehensive intelligence test. For example, the Wonderlic isn't an effective test for identifying or distinguishing gradations among very intelligent, or very deficient, people. For occupations associated with superior intelligence, such as lawyer, physician, engineer, or chemist, intelligence tests having higher ceilings can help distinguish between the brilliant and the merely bright. Examples of such tests are versions of the Stanford-Binet, Woodcock-Johnson, Wechsler, and Ravens Advanced. All of these must be administered by trained psychologists.

The WPT is better when used to create broad categories, such as above average, average, and below average; or "probably qualified" and "probably not qualified". Since the WPT has arguably been taken by more people than any other intelligence test used for pre-employment screening (Wonderlic scores can also be used for admission into Mensa and other high-iq groups), it has amassed considerable predictive validity. The WPT is also inexpensive and easily administered, even online.

Although intelligence tests have demonstrated utility and scientific validity in recruiting and screening potential employees, they must be used lawfully, carefully and with judgement. The most important rule is not to rely on intelligence tests, in and of themselves, in the decision to accept or reject someone. This is, by the way, a good rule to follow when using such tests in other ways, such as admitting people to graduate school, deciding who will be moved into a gifted education program, or who needs cognitive rehabilitation after suffering a stroke.